# Elisabethstraße 59/III, 8010 Graz, Austria

### v1

@@ 1.1 Diplomatics
Sigillography (the study of diplomatic seals), forgery detection, document identification, dating, etc. are all strongly dependent on analysing visual attributes of the documents making image analysis tools indispensable. Although diplomatics data has been part of larger datasets of historical document images, pure image analysis on charters has not been that common. Standing out among recent studies, Boroş et al. [1] demonstrates that Named Entity Recognition directly on images works better than with a two stage pipeline and the work by Leipert et al. [8] which addresses the problem of class imbalance in the context of image segmentation. In Fig. 1 two typical charters, one verso and one recto, can be seen.
@@ 2 Layout Analysis
In order to make sure we gain a proper understanding of the data available in monasterium.net, we randomly chose 1000 charters from it. The heterogeneous nature of the dataset is illustrated by Fig. 2, in which we plot the distribution of width, height, and surface in pixels and mega-pixels, respectively. Additionally, the number of diplomatic seals occurring in each charter image is given.
The dataset not only aims at understanding the quality and nature of the diplomatics data in monasterium.net but also at supporting the creation of pipelines that will offer distant viewing and distant reading. Specifically, we model the layout analysis problem as an object detection problem. As our goal is to maximize the efficiency of diplomatic scholars, we devised an annotation strategy that would require as little time as possible per charter. Diplomatic charters are considered highly formulaic [3], commonly containing a single dominant text-block as can be seen in Fig. 1. Since the images do not seem to need any dewarping, we modeled all layout annotation tasks as defining two-point rectangles, i. e., a rectangle extending from the top left corner to the bottom right corner. Each rectangle belongs to one of 11 specific classes. While these classes are from the point of view of annotation flat and independent, they imply a hierarchical structure:
@@ 2.2 Object Detection
The mean average precision (mAP) for all classes at Intersection over Union (IoU) obtained is , while at IoU it is . These numbers demonstrate very good results, for some classes, the results are even nearly flawless. The confusion matrix is shown in Fig. 3. We can see that classes Img:Seal, Img:WriteableArea, and Img:CalibrationCard all score at precision or higher. These classes are the ones for which annotators only need little interpretation. Class "Wr:OldText" is also performing fairly well with an accuracy of , especially if one takes into account that of missed objects are detected as "Wr:OldNote", which indeed can easily be confused. This class is also quite important as it is the one that detects text blocks, which can subsequently be segmented into text lines so that they can be passed to a handwritten text recognition (HTR) process.
Fig. 4 shows the precision recall curves of each class along with their mAP at IoU. It can be observed that the mAP scores are significantly higher than the precision values reported in the confusion matrix.
@@ 2.3 Annotations with FRAT
We introduce a lightweight annotation tool for document image annotation. The tool can annotate rectangles by just two-points with an arbitrary set of classes. Other than a class, each rectangle can also have an associated Unicode transcription as well as a comment. The tool is called Fast Rectangle Annotation Tool (FRAT). It is web-based and written in Python and Javascript and therefore runs on Linux and Mac OsX natively, and on Windows using Docker. FRAT is available in pypi and can be easily deployed with pip. In order to accommodate this very generic data model, the tool has its own JSON-based file format reflecting these fundamental assumptions and nothing more. The User Interface (UI) is aimed at being extremely lightweight and fast, to a certain extent at the expense of precision. It is inspired by gaming interfaces, where all buttons are accessible as single keys on the keyboard, see Fig. 5. Set operations such as rectangle union and subtraction allow the annotator to figure out efficient use patterns that best suit the domain of data employed.
@@ 3.1 Baseline
For a baseline, we considered the calibration-card object detector described in Section 2. Specifically, since almost all calibration cards are variants of the Kodak calibration cards, they have a length of , even if they differ in other details, while having a height of less than . The calibration cards occur always on the sides of the images vertically or horizontally aligned either vertically or horizontally, by choosing the largest side of the bounding box, we can with high certainty obtain an estimate on how many pixels make a length of 20.5 cm on the photographed plane. This heuristic interpretation of the calibration card object detection gave a Spearman’s correlation of 89.13% and makes it a reasonably strong baseline. To the knowledge of the authors there is no other baseline available in the literature that is applicable in this experimental context.
@@ 3.2 ResResNet
We introduce a network called ResResNet which is a typical ResNet [4] where the final fully connected layer is substituted with a small multi-layer perceptron (MLP) consisting of three fully connected layers with 512, 256, and 1 nodes. While the ResResNet can utilise any ResNet architecture as a backend, we mostly experimented with ResNet18 in order to limit the capacity of the model to suit our relatively small dataset. Due to their average pooling layer, ResResNet can operate on images of arbitrary size. Image resolution can be considered as a texture-base analysis. Thus, during inference, we run ResResNet on a sliding window over the initial image and consider the median prediction as a result for the image. Note: in Fig. 7 this inference method is called stable.
@@ 3.4 Results and Analysis
Resolution estimation is a regression task. Hence, typical performance metrics such as Accuracy and F-score are not applicable. For our performance analysis, we use MSE and and we consider Spearmann’s correlation to be more appropriate as the errors do not follow a Gaussian distribution in a very loose sense. MSE quantifies the size of mistakes in total while Spearman’s correlation quantifies how often predictions are correct. Figure 7 shows the results. While both validation and train sets were curated such that they would only refer to images that contained a calibration card out of necessity, they fail to demonstrate the base limitation of the baseline system, which is that a calibration card is required for an estimate to be possible at all.
Therefore, we gradually remove samples from the validation set from the most erroneous one (according to the squared error) to the least erroneous one. The results are shown in Fig. 8. We notice that the calibration card baseline performs the best for most samples but overall ResResNet is more reliable as it handles hard samples more gracefully. Furthermore, ResResNet deals with images that do not contain any calibration card at all. The fact that “texture” analysis is enough to learn to predict physical resolution indicates that historical textual data is by nature scale-sensitive and because the trained image analysis models operate on various scales, we can speculate that they implicitly learn multiple scale representations of the same phenomenon. Further experiments that might help validate these hypotheses would be to see whether re-scaling train and test sets for typical textural image analysis tasks such as binarization would improve performance on these tasks. Another approach to better make use of this hypothesis would be to add ResResNet as a Spatial Transformer Network (STN) [5] with PpCm being the single degree of freedom.

### v2

@@ 1.1 Diplomatics
Sigillography (the study of diplomatic seals), forgery detection, document identification, dating, etc. are all strongly dependent on analysing visual attributes of the documents making image analysis tools indispensable. Although diplomatics data has been part of larger datasets of historical document images, pure image analysis on charters has not been that common. Standing out among recent studies, Boroş et al. [1] demonstrates that Named Entity Recognition directly on images works better than with a two stage pipeline and the work by Leipert et al. [8] which addresses the problem of class imbalance in the context of image segmentation. In Fig.˜1 two typical charters, one verso and one recto, can be seen.
@@ 2 Layout Analysis
In order to make sure we gain a proper understanding of the data available in monasterium.net, we randomly chose 1000 charters from it. The heterogeneous nature of the dataset is illustrated by Fig.˜2, in which we plot the distribution of width, height, and surface in pixels and mega-pixels, respectively. Additionally, the number of diplomatic seals occurring in each charter image is given.
The dataset not only aims at understanding the quality and nature of the diplomatics data in monasterium.net but also at supporting the creation of pipelines that will offer distant viewing and distant reading. Specifically, we model the layout analysis problem as an object detection problem. As our goal is to maximize the efficiency of diplomatic scholars, we devised an annotation strategy that would require as little time as possible per charter. Diplomatic charters are considered highly formulaic [3], commonly containing a single dominant text-block as can be seen in Fig.˜1. Since the images do not seem to need any dewarping, we modeled all layout annotation tasks as defining two-point rectangles, i. e., a rectangle extending from the top left corner to the bottom right corner. Each rectangle belongs to one of 11 specific classes. While these classes are from the point of view of annotation flat and independent, they imply a hierarchical structure:
@@ 2.2 Object Detection
The mean average precision (mAP) for all classes at Intersection over Union (IoU) obtained is , while at IoU it is . These numbers demonstrate very good results, for some classes, the results are even nearly flawless. The confusion matrix is shown in Fig.˜3. We can see that classes Img:Seal, Img:WriteableArea, and Img:CalibrationCard all score at precision or higher. These classes are the ones for which annotators only need little interpretation. Class "Wr:OldText" is also performing fairly well with an accuracy of , especially if one takes into account that of missed objects are detected as "Wr:OldNote", which indeed can easily be confused. This class is also quite important as it is the one that detects text blocks, which can subsequently be segmented into text lines so that they can be passed to a handwritten text recognition (HTR) process.
Fig.˜4 shows the precision recall curves of each class along with their mAP at IoU. It can be observed that the mAP scores are significantly higher than the precision values reported in the confusion matrix.
@@ 2.3 Annotations with FRAT
We introduce a lightweight annotation tool for document image annotation. The tool can annotate rectangles by just two-points with an arbitrary set of classes. Other than a class, each rectangle can also have an associated Unicode transcription as well as a comment. The tool is called Fast Rectangle Annotation Tool (FRAT). It is web-based and written in Python and Javascript and therefore runs on Linux and Mac OsX natively, and on Windows using Docker. FRAT is available in pypi and can be easily deployed with pip. In order to accommodate this very generic data model, the tool has its own JSON-based file format reflecting these fundamental assumptions and nothing more. The User Interface (UI) is aimed at being extremely lightweight and fast, to a certain extent at the expense of precision. It is inspired by gaming interfaces, where all buttons are accessible as single keys on the keyboard, see Fig.˜5. Set operations such as rectangle union and subtraction allow the annotator to figure out efficient use patterns that best suit the domain of data employed.
@@ 3.1 Baseline
For a baseline, we considered the calibration-card object detector described in Section˜2. Specifically, since almost all calibration cards are variants of the Kodak calibration cards, they have a length of , even if they differ in other details, while having a height of less than . The calibration cards occur always on the sides of the images vertically or horizontally aligned either vertically or horizontally, by choosing the largest side of the bounding box, we can with high certainty obtain an estimate on how many pixels make a length of 20.5 cm on the photographed plane. This heuristic interpretation of the calibration card object detection gave a Spearman’s correlation of 89.13% and makes it a reasonably strong baseline. To the knowledge of the authors there is no other baseline available in the literature that is applicable in this experimental context.
@@ 3.2 ResResNet
We introduce a network called ResResNet which is a typical ResNet [4] where the final fully connected layer is substituted with a small multi-layer perceptron (MLP) consisting of three fully connected layers with 512, 256, and 1 nodes. While the ResResNet can utilise any ResNet architecture as a backend, we mostly experimented with ResNet18 in order to limit the capacity of the model to suit our relatively small dataset. Due to their average pooling layer, ResResNet can operate on images of arbitrary size. Image resolution can be considered as a texture-base analysis. Thus, during inference, we run ResResNet on a sliding window over the initial image and consider the median prediction as a result for the image. Note: in Fig.˜7 this inference method is called stable.
@@ 3.4 Results and Analysis
Resolution estimation is a regression task. Hence, typical performance metrics such as Accuracy and F-score are not applicable. For our performance analysis, we use MSE and and we consider Spearmann’s correlation to be more appropriate as the errors do not follow a Gaussian distribution in a very loose sense. MSE quantifies the size of mistakes in total while Spearman’s correlation quantifies how often predictions are correct. Figure˜7 shows the results. While both validation and train sets were curated such that they would only refer to images that contained a calibration card out of necessity, they fail to demonstrate the base limitation of the baseline system, which is that a calibration card is required for an estimate to be possible at all.
Therefore, we gradually remove samples from the validation set from the most erroneous one (according to the squared error) to the least erroneous one. The results are shown in Fig.˜8. We notice that the calibration card baseline performs the best for most samples but overall ResResNet is more reliable as it handles hard samples more gracefully. Furthermore, ResResNet deals with images that do not contain any calibration card at all. The fact that “texture” analysis is enough to learn to predict physical resolution indicates that historical textual data is by nature scale-sensitive and because the trained image analysis models operate on various scales, we can speculate that they implicitly learn multiple scale representations of the same phenomenon. Further experiments that might help validate these hypotheses would be to see whether re-scaling train and test sets for typical textural image analysis tasks such as binarization would improve performance on these tasks. Another approach to better make use of this hypothesis would be to add ResResNet as a Spatial Transformer Network (STN) [5] with PpCm being the single degree of freedom.